# SpeeDe3DGS: Speedy Deformable 3D Gaussian Splatting with Temporal Pruning and Motion Grouping

### v3

@@ A.3 Ablation on Number of Groups
We ablate the number of motion groups for GroupFlow on the NeRF-DS [yan2023nerf] and D-NeRF [pumarola2020dnerf] datasets using the DeformableGS [yang2023deformable3dgs] codebase. Table 10 reports quantitative results averaged over three independent runs to reduce training variance. All FPS and training time measurements are collected on RTX 3090 and RTX A5000 GPUs.
@@
| NeRF-DS [yan2023nerf] | D-NeRF [pumarola2020dnerf] | ||||||||||||||
@@
A.4 Ablation on Pruning Percentages
Figure 5 illustrates parameter sweeps over soft (densification-stage) and hard (post-densification) pruning percentages for the NeRF-DS [yan2023nerf] and D-NeRF [pumarola2020dnerf] datasets using the SpeeDe3DGS codebase. Each configuration is evaluated in intervals and repeated three times to reduce variance, with results averaged across all runs. All experiments are performed without Temporal Sensitivity Sampling (TSS) or GroupFlow to isolate the effect of Temporal Sensitivity Pruning (TSP) alone. FPS and training time are collected on a RTX A5000 GPU. We empirically select the (, ) soft–hard pruning ratio as it provides the best balance between rendering speed and visual quality.
A.5 Per-Scene Metrics for SpeeDe3DGS
Tables 11, 12, and 13 report per-scene PSNR, SSIM, and LPIPS, respectively, using our standalone SpeeDe3DGS codebase on NeRF-DS [yan2023nerf] and D-NeRF [pumarola2020dnerf]. Tables 14 and 15 provide per-scene rendering FPS and training time in seconds for NeRF-DS, D-NeRF, and the eight HyperNeRF [park2021hypernerf] scenes used in DeformableGS [yang2023deformable3dgs]. As in Appendix A.2, rendering FPS is measured on RTX 3090 GPUs, while training time is measured on RTX A5000 GPUs for NeRF-DS and D-NeRF and on RTX A6000 GPUs for HyperNeRF.
| NeRF-DS [yan2023nerf] | D-NeRF [pumarola2020dnerf] | ||||||||||||||||
@@ A.5 Per-Scene Metrics for SpeeDe3DGS
| NeRF-DS [yan2023nerf] | D-NeRF [pumarola2020dnerf] | ||||||||||||||||
@@
| NeRF-DS [yan2023nerf] | D-NeRF [pumarola2020dnerf] | ||||||||||||||||
@@
| NeRF-DS [yan2023nerf] | D-NeRF [pumarola2020dnerf] | HyperNeRF [park2021hypernerf] | |||||||||||||||||||||||
@@
| NeRF-DS [yan2023nerf] | D-NeRF [pumarola2020dnerf] | HyperNeRF [park2021hypernerf] | |||||||||||||||||||||||
@@
A.6 Qualitative Results on MonoDyGauBench
Figure 6 presents qualitative comparisons between the DeformableGS [yang2023deformable3dgs] baseline and our SpeeDe3DGS methods, using representative scenes from HyperNeRF [park2021hypernerf], Nerfies [park2021nerfies], and iPhone [guo2022neural] within the MonoDyGauBench framework [liang2025monocular]. Pruning (TSP+TSS) and GroupFlow significantly accelerate rendering while maintaining comparable or better visual fidelity than the baseline. As discussed in Appendix A.1, grouped SE(3) motion distillation also acts as a regularizer in real-world scenes – enhancing temporal coherence, mitigating drift under noisy or unstable camera trajectories, and improving overall image quality. This effect is especially pronounced on the creeper scene from iPhone, where the SSIM of the Pruning+GroupFlow model is twice that of both the baseline and Pruning-only variants. Note that our standalone SpeeDe3DGS codebase achieves superior rendering speed and visual fidelity compared to the standardized MonoDyGauBench framework.

### v4

@@ A.3 Ablation on Number of Groups
We ablate the number of motion groups for GroupFlow on the NeRF-DS [52] and D-NeRF [39] datasets using the DeformableGS [54] codebase. Table 10 reports quantitative results averaged over three independent runs to reduce training variance. All FPS and training time measurements are collected on RTX 3090 and RTX A5000 GPUs.
@@
| NeRF-DS [52] | D-NeRF [39] | ||||||||||||||
@@
A.4 Ablation on TSS Hyperparameters
Table 11 ablates the perturbation magnitude and annealing period for Temporal Sensitivity Sampling (TSS) on the NeRF-DS [52] dataset using the SpeeDe3DGS codebase. TSS improves pruning performance across all configurations, and its hyperparameters are not sensitive because the changes between timesteps in real-time video are small. We adopt the same and as DeformableGS [54].
| Method | PSNR | SSIM | LPIPS | ||
| TSP | - | - | 23.78 | 0.8507 | 0.1863 |
| TSP + TSS | 0.05 | 20K | 23.82 | 0.8513 | 0.1861 |
| TSP + TSS | 0.1 | 15K | 23.82 | 0.8510 | 0.1858 |
| TSP + TSS | 0.1 | 20K | 23.81 | 0.8515 | 0.1853 |
| TSP + TSS | 0.1 | 25K | 23.80 | 0.8517 | 0.1853 |
| TSP + TSS | 0.2 | 20K | 23.81 | 0.8514 | 0.1858 |
A.5 Ablation on Pruning Percentages
Figure 5 illustrates parameter sweeps over soft (densification-stage) and hard (post-densification) pruning percentages for the NeRF-DS [52] and D-NeRF [39] datasets using the SpeeDe3DGS codebase. Each configuration is evaluated in intervals and repeated three times to reduce variance, with results averaged across all runs. All experiments are performed without Temporal Sensitivity Sampling (TSS) or GroupFlow to isolate the effect of Temporal Sensitivity Pruning (TSP) alone. FPS and training time are collected on an RTX A5000 GPU. We empirically select the (, ) soft–hard pruning ratio as it provides the best balance between rendering speed and visual quality.
A.6 Per-Scene Metrics for SpeeDe3DGS
Tables 12, 13, and 14 report per-scene PSNR, SSIM, and LPIPS, respectively, using our standalone SpeeDe3DGS codebase on NeRF-DS [52] and D-NeRF [39]. Tables 15 and 16 provide per-scene rendering FPS and training time in seconds for NeRF-DS, D-NeRF, and the eight HyperNeRF [37] scenes used in DeformableGS [54]. As in Appendix A.2, rendering FPS is measured on an RTX 3090 GPU, while training time is measured on an RTX A5000 GPU for NeRF-DS and D-NeRF and on an RTX A6000 GPU for HyperNeRF.
| NeRF-DS [52] | D-NeRF [39] | ||||||||||||||||
@@ A.6 Per-Scene Metrics for SpeeDe3DGS
| NeRF-DS [52] | D-NeRF [39] | ||||||||||||||||
@@
| NeRF-DS [52] | D-NeRF [39] | ||||||||||||||||
@@
| NeRF-DS [52] | D-NeRF [39] | HyperNeRF [37] | |||||||||||||||||||||||
@@
| NeRF-DS [52] | D-NeRF [39] | HyperNeRF [37] | |||||||||||||||||||||||
@@
A.7 Qualitative Results on MonoDyGauBench
Figure 6 presents qualitative comparisons between the DeformableGS [54] baseline and our SpeeDe3DGS methods, using representative scenes from HyperNeRF [37], Nerfies [36], and iPhone [10] within the MonoDyGauBench framework [25]. Pruning (TSP+TSS) and GroupFlow significantly accelerate rendering while maintaining comparable or better visual fidelity than the baseline. As discussed in Appendix A.1, grouped SE(3) motion distillation also acts as a regularizer in real-world scenes – enhancing temporal coherence, mitigating drift under noisy or unstable camera trajectories, and improving overall image quality. This effect is especially pronounced on the creeper scene from iPhone, where the SSIM of the Pruning+GroupFlow model is twice that of both the baseline and Pruning-only variants. Note that our standalone SpeeDe3DGS codebase achieves superior rendering speed and visual fidelity compared to the standardized MonoDyGauBench framework.
A.8 Baseline and Motion Model Limitations
Our methods inherit the limitations of the dynamic Gaussian Splatting frameworks into which they are integrated. In particular, highly deformable scenes remain challenging for current approaches. MonoDyGauBench [25] finds that the baseline methods struggle with non-rigid motion and noisy camera poses, so our approach faces similar limitations.
Nevertheless, Sections 5.2 and Appendices A.1 and A.3 demonstrate that GroupFlow can introduce regularizing effects that improve temporal stability in such scenarios. In practice, globally non-rigid scenes consist of clusters of small Gaussians whose motion is locally rigid due to cloning and splitting during densification. Increasing the number of groups enables progressively finer motion modeling; in the extreme case , GroupFlow distills per-Gaussian neural motion into efficient SE(3) transformations. Adaptive strategies that refine groups based on motion variation could further improve modeling of complex non-rigid motion.